# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v1

@@ 1.5.3 Conversion from KEGG ID to MeSH ID
Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
Top1 Y1 Random 0.000 0.000 0.003 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.250 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
Y2 Random 0.005 0.005 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.500 0.481 0.350 0.318 0.305 0.299 0.328 0.326 0.318 0.376 0.318
Top10 Y1 Random 0.008 0.010 0.008 0.002 0.003 0.001 0.000 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.500 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
Y2 Random 0.041 0.022 0.007 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.727 0.785 0.699 0.657 0.665 0.679 0.735 0.744 0.702 0.773 0.764
MRR Y1 Random 0.012 0.007 0.005 0.003 0.002 0.002 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.331 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
Y2 Random 0.017 0.010 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.590 0.592 0.476 0.434 0.425 0.429 0.468 0.463 0.451 0.513 0.469
@@
1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$\sum_{p\in\mathcal{P}}\left|D_{p}^{y\mid U_{y}}\right|$ 14 104 314 507 521 601 526 439 312 201 67
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}|$ 16 115 363 572 573 732 644 529 401 264 98
$\sum_{p\in\mathcal{P}}|D_{p}^{y}|$ 26 156 513 959 1363 1728 1999 2205 2419 2711 2821
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y}|$ 30 174 566 1028 1419 1794 2068 2263 2480 2773 2886
$\sum_{p\in\mathcal{P}}\left|G_{p}^{y\mid U_{y}}\right|$ 6 29 88 129 195 305 326 269 228 168 62
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}|$ 15 50 177 281 520 935 1087 859 774 682 275
$\sum_{p\in\mathcal{P}}|G_{p}^{y}|$ 6 38 110 170 281 484 658 772 925 1066 1124
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y}|$ 18 68 207 367 652 1356 2015 2359 3087 3518 3627
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{\left|[d]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 4.000 2.500 5.921 8.148 11.407 12.667 14.319 14.649 14.631 15.465 15.977
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}}\left\{|[d]_{p}^{y\mid U_{y}}|\right\}\right\}$ 2.333 4.160 6.038 7.243 6.432 7.805 6.188 5.776 5.032 4.467 3.190
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid L_{y}}\right|\right\}\right\}$ 1.000 1.136 1.789 1.918 2.444 3.441 4.483 5.216 5.529 6.171 6.400
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}}\left\{\left|[g]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 1.000 1.160 1.692 1.843 2.407 3.961 3.835 3.539 3.677 3.733 2.952
Table S9: Statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2.
@@ 2.2 Analogy tasks for drug-gene pairs by year
Metric Setting Method 1970 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
Top1 P1Y1 Random 0.014 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.653 0.515 0.484 0.394 0.303 0.275 0.263 0.256 0.285 0.269
P1Y2 Random 0.000 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.780 0.674 0.656 0.651 0.633 0.636 0.627 0.630 0.659 0.623
P2Y1 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.625 0.467 0.437 0.358 0.250 0.226 0.221 0.209 0.236 0.184
P2Y2 Random 0.005 0.003 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.744 0.633 0.620 0.627 0.613 0.616 0.619 0.621 0.653 0.617
Top10 P1Y1 Random 0.029 0.020 0.009 0.003 0.003 0.001 0.002 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.800 0.670 0.727 0.638 0.597 0.573 0.627 0.566 0.665 0.612
P1Y2 Random 0.026 0.012 0.005 0.004 0.002 0.002 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.906 0.842 0.863 0.863 0.879 0.896 0.913 0.899 0.925 0.911
P2Y1 Random 0.000 0.013 0.006 0.002 0.002 0.001 0.001 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.762 0.627 0.674 0.590 0.533 0.502 0.567 0.513 0.563 0.480
P2Y2 Random 0.032 0.013 0.007 0.004 0.002 0.002 0.001 0.001 0.000 0.000 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.872 0.814 0.832 0.845 0.866 0.883 0.909 0.896 0.922 0.912
MRR P1Y1 Random 0.032 0.008 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.727 0.583 0.582 0.483 0.404 0.378 0.392 0.373 0.410 0.391
P1Y2 Random 0.023 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.843 0.749 0.743 0.731 0.722 0.734 0.737 0.732 0.758 0.731
P2Y1 Random 0.012 0.008 0.006 0.003 0.002 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.695 0.532 0.529 0.444 0.345 0.323 0.341 0.318 0.340 0.282
P2Y2 Random 0.019 0.010 0.006 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.808 0.711 0.708 0.709 0.704 0.716 0.729 0.725 0.752 0.727
Table S10: Gene prediction performance in settings P1Y1, P1Y2, P2Y1, and P2Y2.
@@ 2.2.1 Global setting by year
Table S7 shows the statistics for settings Y1 and Y2. As seen in Table S2, in setting Y1, the query is $d\in D^{y\mid U_{y}}$, and the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. In setting Y2, the query is $d\in D^{y}$, and the search space is $\mathcal{G}^{y}$. Table S7 shows that $D^{y\mid U_{y}}$ has increased since 1970 and started to decrease after 1995. On the other hand, $D^{y}$ and $\mathcal{G}^{y}$ show a monotonic increase over the years.
@@ 2.2.2 Pahtway-wise setting by year
Table S9 shows the statistics for settings P1Y1, P1Y2, P2Y1, and P2Y2. As seen in Table S2, the query is $d\in D_{p}^{y\mid U_{y}}$ in setting P1Y1, $d\in D_{p}^{y}$ in setting P1Y2, $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ in P2Y1, and $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ in setting P2Y2. Table S9 shows that $D_{p}^{y\mid U_{y}}$ and $\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ have increased since 1970 and started to decrease after 1995. On the other hand, $D_{p}^{y}$ and $\mathcal{D}_{p}^{y}\cap D^{y}$ show a monotonic increase over the years.
We showed the plots of the results for settings P1Y1 and P1Y2 in Fig. 3. Fig. S2 shows the plots of the results for settings P2Y1 and P2Y2. For settings P1Y1, P1Y2, P2Y1, and P2Y2, Table S10 shows detailed results used in Fig. 3 and Fig. S2. As shown in Table 3, settings P1 and P2 show roughly the same trends. Similarly, in Table S10 , settings P1Y1 and P2Y1, as well as P1Y2 and P2Y2, each show similar trends.

### v2

@@ 1.5.3 Conversion from KEGG ID to MeSH ID
Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
Top1 Y1 Random 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y}$ 0.365 0.262 0.249 0.236 0.247 0.272 0.275 0.287 0.323 0.301 0.300
Y2 Random 0.000 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.333 0.208 0.165 0.132 0.076 0.092 0.062 0.090 0.107 0.021
Top10 Y1 Random 0.016 0.007 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.000
$\hat{\mathbf{v}}^{y}$ 0.837 0.577 0.563 0.562 0.578 0.652 0.691 0.651 0.714 0.693 0.686
Y2 Random 0.015 0.003 0.002 0.003 0.001 0.001 0.000 0.001 0.000 0.000
$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.625 0.487 0.438 0.410 0.347 0.397 0.352 0.344 0.328 0.271
MRR Y1 Random 0.011 0.005 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001 0.000
$\hat{\mathbf{v}}^{y}$ 0.534 0.370 0.353 0.346 0.363 0.399 0.410 0.410 0.455 0.430 0.426
Y2 Random 0.008 0.006 0.003 0.002 0.001 0.001 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}^{y\mid L_{y}}$ 0.438 0.304 0.253 0.214 0.163 0.182 0.147 0.173 0.188 0.103
@@
1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
$\sum_{p\in\mathcal{P}}|D_{p}^{y}|$ 85 536 1092 1668 2271 2609 2897 3122 3447 3578 3612
$\sum_{p\in\mathcal{P}}|G_{p}^{y}|$ 15 76 127 234 444 634 765 916 1088 1167 1178
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y}|$ 397 1015 1559 2055 2394 2690 2909 3132 3458 3581 3614
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y}|$ 25 136 271 553 1242 1896 2325 2570 2952 3174 3186
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in D_{p}^{y}}\{|[d]_{p}^{y}|\}\}$ 1.078 1.446 1.289 1.341 1.603 1.803 1.841 2.027 1.983 1.979 1.990
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in G_{p}^{y}}\{|[g]_{p}^{y}|\}\}$ 6.143 8.485 8.084 7.485 7.027 6.032 5.607 5.529 4.973 4.830 4.847
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{d\in\mathcal{D}_{p}^{y}\cap D^{y}}\{|[d]^{y}|\}\}$ 1.078 1.430 1.356 1.616 2.211 2.532 2.724 2.902 2.749 2.759 2.776
$\mathrm{E}_{p\in\mathcal{P}}\{\mathrm{E}_{g\in\mathcal{G}_{p}^{y}\cap G^{y}}\{|[g]^{y}|\}\}$ 10.000 13.496 12.787 11.633 8.815 7.163 6.968 6.505 6.525 6.692 6.714
Table S13: Statistics for settings P1Y1 and P2Y1.
Year
1975 1980 1985 1990 1995 2000 2005 2010 2015 2020
$\sum_{p\in\mathcal{P}}\left|D_{p}^{y\mid U_{y}}\right|$ 40 253 419 448 565 500 426 303 195 66
$\sum_{p\in\mathcal{P}}\left|G_{p}^{y\mid U_{y}}\right|$ 10 58 90 154 272 295 250 218 163 61
$\sum_{p\in\mathcal{P}}|\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}|$ 42 275 464 492 681 612 511 387 257 96
$\sum_{p\in\mathcal{P}}|\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}|$ 14 68 121 250 531 607 589 549 496 206
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in D_{p}^{y\mid U_{y}}}\left\{\left|[d]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 1.034 1.445 1.247 1.285 1.477 1.571 1.596 1.724 1.648 1.614
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in G_{p}^{y\mid U_{y}}}\left\{\left|[g]_{p}^{y\mid U_{y}}\right|\right\}\right\}$ 4.194 7.157 6.560 5.761 5.107 4.329 3.924 3.884 3.538 3.443
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}}\left\{\left|[d]^{y\mid U_{y}}\right|\right\}\right\}$ 1.034 1.427 1.301 1.497 1.925 2.126 2.288 2.376 2.176 2.113
$\mathrm{E}_{p\in\mathcal{P}}\left\{\mathrm{E}_{g\in\mathcal{G}_{p}^{y}\cap G^{y\mid U_{y}}}\left\{\left|[g]^{y\mid U_{y}}\right|\right\}\right\}$ 7.000 10.525 9.027 7.890 5.667 4.729 4.392 4.155 4.096 4.249
Table S14: Statistics for settings P1Y2 and P2Y2.
1.6 Details of predictions using the GPT series
In this study, we used GPT-3.5, GPT-4, and GPT-4o as baselines for generative models. The details of the models are shown in Table S5. The prompts are shown in Table S6.
@@ 2.1.2 Prediction of drugs from genes
2.1.3 Predictions by GPT models
The prediction results from the GPT models are shown in Table S9. The set of query drugs and the set of correct target genes depend not only on the settings G, P1, and P2 but also on the vocabulary set. Here, we conducted experiments using vocabularies obtained from BioConceptVec or our skip-gram embeddings. The best accuracies for all combinations of vocabularies and settings are summarized in the Table 3 in the main text.
@@ 2.2 Analogy tasks for drug-gene pairs by year
Metric Setting Method 1975 1980 1985 1990 1995 2000 2005 2010 2015 2020 2023
Top1 P1Y1 Random 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 0.706 0.621 0.613 0.592 0.589 0.588 0.588 0.587 0.614 0.582 0.589
P2Y1 Random 0.002 0.001 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 0.638 0.562 0.568 0.570 0.581 0.581 0.596 0.597 0.626 0.591 0.600
P1Y2 Random 0.003 0.001 0.001 0.000 0.000 0.001 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.350 0.482 0.442 0.366 0.290 0.266 0.254 0.244 0.272 0.258
P2Y2 Random 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.333 0.444 0.399 0.333 0.244 0.221 0.213 0.202 0.226 0.177
Top10 P1Y1 Random 0.014 0.009 0.004 0.002 0.001 0.001 0.001 0.001 0.000 0.001 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 1.000 0.871 0.862 0.835 0.847 0.881 0.889 0.867 0.886 0.869 0.862
P2Y1 Random 0.015 0.006 0.004 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 0.915 0.797 0.812 0.811 0.842 0.877 0.899 0.881 0.900 0.884 0.880
P1Y2 Random 0.018 0.006 0.004 0.002 0.002 0.002 0.001 0.001 0.000 0.002
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.625 0.648 0.704 0.621 0.586 0.564 0.620 0.558 0.662 0.606
P2Y2 Random 0.012 0.007 0.003 0.002 0.001 0.001 0.001 0.001 0.001 0.001
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.595 0.615 0.655 0.575 0.530 0.495 0.562 0.506 0.560 0.479
MRR P1Y1 Random 0.012 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y}$ 0.849 0.724 0.712 0.684 0.680 0.690 0.694 0.688 0.713 0.684 0.685
P2Y1 Random 0.008 0.006 0.003 0.002 0.002 0.001 0.001 0.001 0.001 0.001 0.001
$\hat{\mathbf{v}}_{p}^{y}$ 0.772 0.658 0.664 0.661 0.673 0.684 0.704 0.700 0.726 0.696 0.700
P1Y2 Random 0.008 0.007 0.002 0.002 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.488 0.555 0.547 0.459 0.392 0.370 0.384 0.362 0.400 0.382
P2Y2 Random 0.012 0.005 0.003 0.003 0.001 0.001 0.001 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}^{y\mid L_{y}}$ 0.466 0.514 0.499 0.423 0.340 0.316 0.333 0.312 0.333 0.277
Table S15: Gene prediction performance in settings P1Y1, P2Y1, P1Y2, and P2Y2.
@@ 2.2.1 Global setting by year
Tables S10 and S11 show the statistics for settings Y1 and Y2. As seen in Table S2, in setting Y1, the query is $d\in D^{y}$, and the search space is $\mathcal{G}^{y}$. In setting Y2, the query is $d\in D^{y\mid U_{y}}$, and the search space is $\mathcal{G}^{y}\setminus[d]^{y\mid L_{y}}$. In Table S10, $D^{y}$ and $\mathcal{G}^{y}$ show a monotonic increase over the years. On the other hand, Table S11 shows that $D^{y\mid U_{y}}$ has increased since 1975 and started to decrease after 1995.
@@ 2.2.2 Pahtway-wise setting by year
Tables S13 and S14 show the statistics for settings P1Y1, P2Y1, P1Y2, and P2Y2. As seen in Table S2, the query is $d\in D_{p}^{y}$ in setting P1Y1, $d\in\mathcal{D}_{p}^{y}\cap D^{y}$ in setting P2Y1, $d\in D_{p}^{y\mid U_{y}}$ in setting P1Y2, and $d\in\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ in P2Y2. In Table S13, $D_{p}^{y}$ and $\mathcal{D}_{p}^{y}\cap D^{y}$ show a monotonic increase over the years. On the other hand, Table S14 shows that $D_{p}^{y\mid U_{y}}$ and $\mathcal{D}_{p}^{y}\cap D^{y\mid U_{y}}$ have increased since 1975 and started to decrease after 1995.
We showed the plots of the results for settings P1Y1 and P1Y2 in Fig. 3. Fig. S2 shows the plots of the results for settings P2Y1 and P2Y2. For settings P1Y1, P2Y1, P1Y2, and P2Y2, Table S15 shows detailed results used in Fig. 3 and Fig. S2. As shown in Table 3, settings P1 and P2 show roughly the same trends. Similarly, in Table S15, settings P1Y1 and P2Y1, as well as P1Y2 and P2Y2, each show similar trends.